# Spectral Properties of Generalized Correlation Matrices

### v1

@@ Abstract.
We introduce a family of coefficients that generalize the notion of correlation and explore their properties in the large dimensional multivariate case, showing that in the null case of uncorrelated variables, the spectrum of generalized correlation matrices is distributed according to an affine transformation of the Marčenko-Pastur law.
@@ 1. Introduction
Building on the work of [3], we show that the spectrum of generalized correlation matrices, with entries being generalized correlation coefficients, are also distributed according to an affine transformation of a MP law. This novel result, illustrated here by Figure 1, highlights the intricate relationship between MP distributions and sample correlation matrices of uncorrelated variables.
@@ 2. Main result
be an array of i.i.d. random variables taking values in a measured space $\mathcal{S}$, and $\phi:\mathcal{S}^{d}\to\mathbb{R}$ be an antisymmetric measurable function (all, except $d$, implicitly depending on $n$). We introduce

### v2

# Random Matrices and U-Statistics

@@ Abstract.
We introduce a family of coefficients based on U-statistics that generalize the notion of correlation and explore their properties in the large dimensional multivariate case, showing that in the null case of uncorrelated variables, the spectrum of generalized correlation matrices is distributed according to an affine transformation of the Marčenko-Pastur law.
@@ 1. Introduction
Building on the work of [3], we show that the spectrum of generalized correlation matrices, with entries being U-statistics, are also distributed according to an affine transformation of a MP law. This novel result, illustrated here by Figure 1, highlights the intricate relationship between MP distributions and sample correlation matrices of uncorrelated variables.
@@ 2. Main result
be an array of i.i.d. random variables taking values in a measured space, and $\phi:\mathcal{S}^{d}\to\mathbb{R}$ be an antisymmetric measurable function (all, except $d$, implicitly depending on $n$). We introduce
@@
Remark 2.7.
For any $\mathsf{u},\mathsf{v}\in\mathcal{X}^{n}$ and $d\geq 1$, the $\phi$-covariance $\mathcal{COV}^{\phi}(\mathsf{u},\mathsf{v})$ is in fact a U-statistic of degree $d$ with (symmetric) kernel $h(\mathsf{z}_{1},\ldots,\mathsf{z}_{d})=\phi(\mathsf{u}_{1},\ldots,\mathsf{u}_{d})\phi(\mathsf{v}_{1},\ldots,\mathsf{v}_{d})$ where $\mathsf{z}_{i}=(\mathsf{u}_{i},\mathsf{v}_{i})$. This property infers to $\phi$-covariances many properties like law of large numbers, asymptotic normality, law of iterated logarithms and large deviations [11, 7].